OpenStudy (anonymous):

find the consecutive even integers such as the first times the third is 2 greater than 5 times the second.

OpenStudy (anonymous):

pls answer I really need help

OpenStudy (shubhamsrg):

if they are a-1,a and a+1 then (a-1)(a+1) = 2 + 5(a) just solve for a

OpenStudy (anonymous):

how to solve for the a

OpenStudy (shubhamsrg):

simplify LHS,,,what do you get ? NOTE: LHS MEANS LEFT-HAND-SIDE!! :P

OpenStudy (anonymous):

sorry i dont know how LHS works

OpenStudy (shubhamsrg):

you know how to solve quadratic eqns atleast ?dont you ?

OpenStudy (anonymous):

can you solve it pls

OpenStudy (shubhamsrg):

yes i can,,but i wont.. how about i ask you the same ques? :P

OpenStudy (anonymous):

Let the numbers be 2n,2n+2,2n+4 then according to the given data 2n(2n+4) -2=5(2n+2) 4n^2+8n -2=10n+10 4n^2-2n-12=0 2(2n^2-n-6)=o now apply the quadratic formula \[(-b +/- \sqrt{b^2 - 4ac}) \div 2a\] we get the values of n to be 2,-3/2 taking +ve value of n we get the numbers 4,6,10

OpenStudy (anonymous):

sorry the numbers are 4,6,8

OpenStudy (shubhamsrg):

ohh i never noticed it said "even" ..hmm..sorry then :P

OpenStudy (anonymous):

hahaha.. carl.. pang college na ung sinagut nia ahh.. xD di yan tatanggapin ni sir

OpenStudy (anonymous):

panu kasi derek

OpenStudy (anonymous):

hahaha. magtanung ka ulit sa ibang tao.. i-bump mo ulit ung post

OpenStudy (anonymous):

I NEED SOLUTION FOR A HIGH SCHOOL STUDENT

OpenStudy (anonymous):

pls pls answer